The highest barrier that a projectile can clear is 10.0 m, when the projectile is launched at an angle of 29.0 ° above the horizontal. What is the projectile's launch speed?


1
Expert's answer
2021-01-27T15:26:40-0500

Let's first find the "y"-component of the initial velocity of the projectile:


"v_y^2=v_{0y}^2-2gy,""0=v_{0y}^2-2gy,""v_{0y}=\sqrt{2gy}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot10.0\ m}=14\ \dfrac{m}{s}."

Finally, we can find the projectile's launch speed from the trigonometry:


"v_0=\dfrac{v_{0y}}{sin\theta}=\dfrac{14\ \dfrac{m}{s}}{sin29.0^{\circ}}=28.87\ \dfrac{m}{s}."